# Geometric Design and Gait Co-Optimization for Soft Continuum Robots Swimming at Low and High Reynolds Numbers

### v1

@@ I Introduction
Swimming has attracted significant attention across various fields, including fluid mechanics [1, 2, 3, 4, 5, 6, 7], biology [8, 9], and robotics [10, 11, 12, 13]. The study of the mechanisms behind the undulatory self-propulsion of aquatic organisms has provided valuable insights [14, 15, 16, 7], inspiring a wide range of research efforts focused on the design and control of robotic swimmers [17, 18, 19, 20, 21].
@@ II Continuum Kinematics
In this section, we derive the kinematics of the soft continuum swimmer using Lie group theory. We begin by assuming that the deformation of the continuum satisfies the area conservation condition. Under this assumption, the deformation is entirely determined by the change in curvature along the backbone. At the normalized time $t$, the transformation from the tail to the frame at the normalized arclength $s$ is given by
@@ II-A Curvature Parameterization
Figure 2: Demonstration of the deformation of soft continuum robots parameterized by curvature functions. (Top left): The curvature of a two-shape-mode swimmer, where the curvature is a sum of two shape modes. The shape modes are visualized by blue and yellow curves at $t=1$, with shape variables represented by dashed lines at $s=0$. (Top right): Snapshots of the two-shape-mode swimmer’s configuration corresponding to the black curves in the top left subfigure, with curvature indicated by a red-white-black color scheme. Each shape mode is also visualized as a solid blue and yellow swimmer on the sides. (Bottom left): The curvature of an infinitely flexible swimmer, parameterized by unit shape modes distributed along the arclength. The unit shape modes, scaled by the shape variables, are visualized by the red-white-black curves. (Bottom right): Snapshots of the infinitely flexible swimmer’s configuration corresponding to the black curves in the bottom left subfigure.
@@ III Continuum Swimming Dynamics
At very high Re, the dynamics are dominated by inertial effects, and viscous drag becomes negligible. In this case, because there are no external forces or torques acting on the combined robot-fluid system, the equations of motion can be derived from the conservation of momentum of the combined robot-fluid system, with the local metrics representing the fluid-added inertia. By considering the soft continuum swimmer as a slender body that is neutrally buoyant, with a circular, constant cross-section along its backbone, we can use Lighthill’s slender body theory to compute the fluid-added inertia matrix of each infinitesimal segment of the continuum [3, 45, 18, 21]:
@@
where the second term represents the fluid-added mass of the circular cross-section, $\rho$ is the fluid density, and $R$ is the cross-sectional radius of the swimmer.
The pulled-back metrics $M$ can be partitioned as
@@ IV-A Gradient of Displacement Due to Periodic Deformation
Figure 4: Comparison of the optimal design-and-gait efficiencies of different continuum swimmers at low and high Re. In both plots, efficiency is normalized by the maximum efficiency. The low Re swimmers use dissipated power metrics based on the Riemannian metrics of the system’s shape velocity, while the high Re swimmer uses the covariant acceleration metric related to the square of the active force, accounting for dynamics.
@@
As discussed in § II-B, the swimmer’s body velocity can be determined by applying the local connection $\mathbf{A}$ to the time derivative of the curvature and then integrating the contributions at each point along the backbone. Consequently, the displacement resulting from a periodic deformation, $\phi$, can be expressed as the line integral of the local connection along the shape change over the deformation period. Building on previous work [51, 17, 35, 20, 50], the exponential coordinates of this displacement, $\log\left(g_{\phi}\right)$, can be approximated by the surface integral of the total Lie bracket $D(\mathbf{A})$ of the local connections: log(g_ϕ) &= ∮_ϕg(t)g(t)dt
= ∫^1_0(∮_ϕg(t)A(t, s)dκ(s))ds
≈∫^1_0∫^1_s(∯_ϕ D(A)_s, σ)dσds, where
@@ V Numerical Results
In this section, we present simulation results for the optimal design and gait of soft continuum swimmers at low and high Re, obtained using the proposed framework. We compare our optimal solutions with three-link and serpenoid swimmers to demonstrate improvements in efficiency. Furthermore, we show that our solution matches or exceeds the efficiency of infinitely flexible swimmers at low Re and swimmers with higher degrees of freedom at high Re.
@@
In our implementation, we discretize the continuum body into 100 segments and use trapezoidal integration for numerical analysis. Each shape mode and shape variable is defined by 10 control points. The infinitely flexible swimmer is parameterized by 10 unit shape modes. At high Re, the swimmer has a slenderness ratio of 100:1, and at low Re, the drag coefficient ratio is 2:1. Simulations are performed using the ODE45 function in MATLAB. Deformation optimization is carried out with the fmincon function in MATLAB, with the gradient input calculated using our method. The optimization typically takes less than 5 minutes.
@@
Figure 5: Comparison of the optimal deformation determined by the design and gait of the two-shape-mode swimmer with other swimmers at low and high Re. (Left): Comparison of the curvature between the optimal design and gait of the two-shape-mode swimmer and the optimal gait of the infinitely flexible swimmer at low Re. (Right): Comparison of the curvature between the optimal design and gait of the two-shape-mode swimmer and the three-shape-mode swimmer at high Re.
Figure 6: Snapshots of the optimal shape mode design and gaits for soft continuum swimmers at low and high Re at different times in the period $T$. The color of the swimmer indicates the ratio of the two shape modes dominating the deformation at each moment. The gray fins represent the drag force or negative linear momentum on each segment of the swimmer at that moment.
@@
Fig. 4 illustrates the optimal design-and-gait efficiencies of different continuum swimmers at low and high Re. Overall, the three-link swimmer exhibits lower efficiency, while the two-shape-mode swimmer with optimal design-and-gait demonstrates high efficiency, even when compared to systems with higher degrees of freedom. For swimming at low Re, we consider the dissipated power metrics based on the Riemannian metrics of the system’s shape velocity. In this scenario, the serpenoid swimmer performs similarly to the optimal two-shape-mode swimmer. Notably, the optimal design of the two-shape-mode swimmer achieves an efficiency nearly equivalent to that of an infinitely flexible swimmer. This similar efficiency highlights the utility of considering design-and-gait co-optimization—only two actuators with the optimal design-and-gait are needed to achieve the same efficiency as a swimmer with infinite degrees of freedom.
@@
Fig. 5 compares the optimal design and gait of the two-shape-mode swimmer with other swimmers at low and high Re, respectively. For swimming at low Re, the comparison of curvature between the two-shape-mode swimmer and the infinitely flexible swimmer shows that the two-shape-mode co-optimization successfully reproduces most of the curvature trends of the infinitely flexible swimmer’s optimal gait, with the exception of subtle details at the head and tail that cannot be captured by the weighted sum of the two shape modes. This results in nearly identical efficiency between them, highlighting the practicality of design and gait co-optimization for hardware implementation for soft continuum swimmers.
@@
Fig. 6 shows the optimal deformation of different swimmers at low and high Re. [^7] Please refer to the supplementary video for more details, including demonstrations of swimming in the $y$ and $\theta$ directions. For the three-link and serpenoid swimmers, we fixed the shape mode and optimized only the gait, while for the two-shape-mode swimmer, both shape modes and gaits were optimized. At low Re, the deformation becomes progressively smoother and more consistent from the three-link to the two-shape-mode swimmer, leading to a more uniform drag distribution on the continuum and reduced power dissipation due to viscous drag.
@@ VI Conclusion and Future Work
Future work includes applying our framework to real robots [34], improving soft actuator models [52], and exploring multi-material soft robots [53]. We will also investigate more complex locomotion tasks [54, 55] and explore more complicated hydrodynamics with non-zero momentum effects [56, 57, 58]. Additionally, we plan to extend this framework to jet propulsion systems [59] and refine coordinate optimization to improve displacement approximation [38].

### v2

@@ I Introduction
Swimming has attracted significant attention across various fields, including fluid mechanics [1, 2, 3, 4, 5, 6, 7, 8], biology [9, 10], and robotics [11, 12, 13, 14]. The study of the mechanisms behind the undulatory self-propulsion of aquatic organisms has provided valuable insights [15, 16, 17, 8], inspiring a wide range of research efforts focused on the design and control of robotic swimmers [18, 19, 20, 21, 22].
@@ II Continuum Kinematics
In this section, we derive the kinematics of the soft continuum swimmer using Lie group theory. We begin by assuming that the deformation of the continuum satisfies the area conservation condition. Under this assumption, the deformation is entirely determined by the change in curvature along the backbone [43]. At the normalized time $t$, the transformation from the tail to the frame at the normalized arclength $s$ is given by
@@ II-A Curvature Parameterization
Figure 2: Demonstration of the deformation of soft continuum robots parameterized by curvature functions. (Top left): The curvature of a two-shape-mode swimmer, where the curvature is a sum of two shape modes. The shape modes are visualized by black and grey curves at $t=1$, with shape variables represented by dashed lines at $s=0$. (Top right): Snapshots of the two-shape-mode swimmer’s configuration corresponding to the black curves in the top left subfigure, with curvature indicated by a red-white-black color scheme. Each shape mode is also visualized as a solid black and grey swimmer on the sides. (Bottom left): The curvature of an infinitely flexible swimmer, parameterized by unit shape modes distributed along the arclength. The unit shape modes, scaled by the shape variables, are visualized by the red-white-black curves. (Bottom right): Snapshots of the infinitely flexible swimmer’s configuration corresponding to the black curves in the bottom left subfigure.
@@ III Continuum Swimming Dynamics
At very high Re, inertial effects dominate and viscous drag becomes negligible. In this case, because there are no external forces or torques acting on the combined robot-fluid system, the equations of motion can be derived from the conservation of momentum of the combined robot-fluid system, with the local metrics representing the fluid-added inertia. By considering the soft continuum swimmer as a slender body that is neutrally buoyant, with a circular, constant cross-section along its backbone, we can use Lighthill’s slender body theory to compute the fluid-added inertia matrix of each infinitesimal segment of the continuum [4, 47, 19, 22]:
@@
where $\rho$ is the fluid density and $R$ is the swimmer’s cross-sectional radius. The first term represents the inertia and the second term corresponds to the fluid-added mass of the circular cross-section. This formulation aligns with Lighthill’s theory, which relates reactive forces to the added mass of fluid that acquires momentum through shape deformations.
The pulled-back metrics, $M$, which encode fluid drag dissipation at low Re or fluid-added inertia at high Re, can be partitioned as
@@ IV-A Gradient of Displacement Due to Periodic Deformation
Figure 4: Comparison of the optimal design-and-gait efficiencies of different continuum swimmers swimming forward at low and high Re. In both plots, efficiency is normalized by the maximum efficiency. The low Re swimmers use dissipated power metrics based on the Riemannian metrics of the system’s shape velocity, while the high Re swimmer uses the covariant acceleration metric related to the square of the active force.
@@
As discussed in § II-B, the swimmer’s body velocity can be determined by applying the local connection $\mathbf{A}$ to the time derivative of the curvature and then integrating the contributions at each point along the backbone. Consequently, the displacement resulting from a periodic deformation, $\phi$, can be expressed as the line integral of the local connection along the shape change over the deformation period. Building on previous work [53, 18, 36, 21, 52], the exponential coordinates of this displacement, $\log\left(g_{\phi}\right)$, can be approximated by the surface integral of the total Lie bracket $D(\mathbf{A})$ of the local connections: g_ϕ&= ∮_ϕg(t)g(t)dt
= ∮_ϕ(∫^1_0g(t)A(t, s)K_t(t, s)ds)dt
log(g_ϕ) ≈∫^1_0∫^1_s(∯_ϕ D(A)_s, σ)dσds, where
@@ V Numerical Results
This section presents simulation results for the optimal design and gait of soft continuum swimmers swimming forward at low and high Re using our framework. [^6] Please refer to the supplementary video for more details, including demonstrations of swimming in the $y$ and $\theta$ directions. We compare our optimal solutions with three-link and serpenoid swimmers to demonstrate efficiency improvements. Furthermore, we show that our solution matches or exceeds the efficiency of infinitely flexible swimmers at low Re and swimmers with higher degrees of freedom at high Re.
@@
In our implementation, we discretize the continuum body into 100 segments and use trapezoidal integration for numerical analysis. Each shape mode and shape variable is defined by 10 control points. The infinitely flexible swimmer is parameterized by 10 unit shape modes. At high Re, the swimmer has a slenderness ratio of 100:1, and at low Re, the drag coefficient ratio is 2:1. Simulations are performed using the ODE45 function in MATLAB. Deformation optimization is carried out with the fmincon function in MATLAB, with the gradient input calculated using our method. For three-link and snake cases, the initial guess was set as the central circle in the shape space, and these solutions served as the initial guess for higher-order optimization. Self-collision was avoided through constraints. Because the method is gradient-based, the initial guess influences the converged local minimum; assuming a single traveling wave typically leads to a single wave solution. As [2] suggests, exploring multiple traveling waves may enhance efficiency—a direction for future work. The convergence criteria were based on optimality and changes in decision variables, with optimization typically completing in under 5 minutes.
@@
Figure 5: Comparison of the optimal forward swimming deformation for determined by the design and gait of the two-shape-mode swimmer with other swimmers at low and high Re. (Left): Comparison of the curvature between the optimal design and gait of the two-shape-mode swimmer and the optimal gait of the infinitely flexible swimmer at low Re. (Right): Comparison of the curvature between the optimal design and gait of the two-shape-mode swimmer and the three-shape-mode swimmer at high Re.
@@
Fig. 4 illustrates the optimal design-and-gait efficiencies of different continuum swimmers at low and high Re. Overall, the three-link swimmer exhibits lower efficiency, while the two-shape-mode swimmer with optimal design-and-gait demonstrates high efficiency, even when compared to systems with higher degrees of freedom. For swimming at low Re, we consider the dissipated power metrics based on the Riemannian metrics of the system’s shape velocity. In this scenario, the serpenoid swimmer performs similarly to the optimal two-shape-mode swimmer. Notably, the optimal design of the two-shape-mode swimmer achieves an efficiency nearly equivalent to that of an infinitely flexible swimmer. This similar efficiency indicates that two shape modes encoding appropriate traveling waves, combined with effective gait modulation, are sufficient to convert all joint actuation power into propulsion without incurring additional dissipation from internal cancellation. This result highlights the utility of considering design-and-gait co-optimization—only two actuators with the optimal design-and-gait are needed to achieve the same efficiency as a swimmer with infinite degrees of freedom.
@@
Fig. 5 compares the optimal design and gait of the two-shape-mode swimmer with other swimmers at low and high Re, respectively. For swimming at low Re, the comparison of curvature between the two-shape-mode swimmer and the infinitely flexible swimmer shows that the two-shape-mode co-optimization successfully reproduces most of the curvature trends of the infinitely flexible swimmer’s optimal gait, with the exception of subtle details at the head and tail that cannot be captured by the weighted sum of the two shape modes. The normalized mean squared error between the two-shape-mode swimmer’s curvature and that of the infinitely flexible swimmer is approximately 1.3%. This results in nearly identical efficiency between them, highlighting the practicality of design and gait co-optimization for hardware implementation for soft continuum swimmers.
@@
Figure 6: Snapshots of the optimal shape mode design and gaits for soft continuum swimmers swimming forward at low and high Re at different times in the period $T$. The color of the swimmer indicates the ratio of the two shape modes dominating the deformation at each moment. The gray fins represent the drag force or negative linear momentum on each segment of the swimmer at that moment.
Fig. 6 shows the optimal deformation of different swimmers at low and high Re. For the three-link and serpenoid swimmers, we fixed the shape mode and optimized only the gait, while for the two-shape-mode swimmer, both shape modes and gaits were optimized. At low Re, the deformation becomes progressively smoother and more consistent from the three-link to the two-shape-mode swimmer, leading to a more uniform drag distribution on the continuum and reduced power dissipation due to viscous drag.
@@
Comparing behaviors at low and high Re reveals that high-Re swimmers exhibit greater bending. This difference may arise from the definitions of power consumption: at low Re, power dissipation scales with the square of velocity, whereas at high Re, it is based on the square of covariant acceleration. Although the dynamics at low and high Re are not directly linked, this insight could inform future studies on swimming at intermediate Re.
@@ VI Conclusion and Future Work
Future work includes applying our framework to real robots [54, 35], refining soft actuator models to account for response speed, durability, and fatigue during optimization [43, 55], and exploring multi-material soft robots [56]. We will also investigate more complex locomotion tasks [57, 58] and explore more complicated hydrodynamics with non-zero momentum effects [59, 60, 61]. Additionally, we plan to extend this framework to jet propulsion systems [62] and refine coordinate optimization to improve displacement approximation [39]. Compared to [2], our solution is consistent in that the motion wave propagates from head to tail. However, our model does not exhibit the linearly increasing amplitude observed in [2], possibly due to the omission of vortex shedding and second-order dynamics. Incorporating a dynamic model could better capture these effects and extend our approach to scenarios at intermediate Reynolds numbers, where both fluid drag and added mass are significant [22], although this would complicate the optimization due to time-variant dynamics.